Calculate the characteristic impedance for a line that exhibits an inductance of 4 nH/m and 1.5 pF/m

1
Expert's answer
2022-02-02T23:02:02-0500

Characteristic impedance line




R  is the resistance per unit length=0

L is the inductance per unit length= 4 nH/m

G is the conductance per unit length=0

C is the capacitance  per unit length= 1.5 pF/m

j is the imaginary

Characteristic impedance line= (L/C)1/2= (4*10-9/1.5*10-15)1/2= 2.67*106 ohm